# Majority voting is not good for heaven or hell, with mirrored performance

### v4

@@ 5 Conclusion
The ViSE model is designed to evaluate the effectiveness of individual voting strategies and social decision rules in environments of varying favorability and, more generally, in environments with various gain distributions. In this paper, we show that in the case of Gaussian environments and two-component (a group and individualists) societies, the symmetrized majority rule is effective only under more or less neutral conditions, while in ‘heaven’ or ‘hell’ it is harmful. Furthermore, for environments with symmetric distributions, the performance of this rule is symmetric about the neutrality point. For general distributions with finite mean, the performance is the same for any two opposite generators. This is true both for the class of voting bodies formed by agents with complementary strategies and a complementary voting rule, and for the even more general class of antisymmetric voting bodies. Returning to two-component societies: in some of them, cooperating group members improve their position at the expense of individualists, which is disadvantageous for the whole society.
@@
Acknowledgments We are very grateful to two anonymous referees for their careful reading and especially for suggesting a generalization and the idea of reverse time to obtain a simpler proof of the technical result. In the final version, we use a similar approach in other terms, supplemented by the condition of antisymmetry of the voting body, which is necessary to generalize the result. We also thank the participants of the seminar on stochastic voting models at the Institute of Control Sciences of the Russian Academy of Sciences: Yana Tsodikova, Anton Loginov, Vitaly Malyshev, Vladislav Maksimov, Anna Khmelnitskaya, Irina Kalushina, Zoya Lezina, Vladimir Borzenko.

### v5

@@ 6 Conclusion
Figure 4: The logical structure of the symmetry and complementarity analysis: an antisymmetric voting body, its properties, and typical elements. Lemmas and corollaries connect the boxes through step-up implications.
The ViSE model is designed to evaluate the effectiveness of individual voting strategies and social decision rules in environments of varying favorability and, more generally, in environments with various gain distributions. In this paper, we show that, in Gaussian environments and two-component (a group and individualists) societies, the symmetrized majority rule is effective only under more or less neutral conditions, while in “heaven” or “hell” it is harmful. Furthermore, for environments with symmetric distributions, the performance of this rule is symmetric about the neutrality point. For general distributions with finite mean, this performance, measured relative to the baseline rule, is the same for any two opposite generators. This is true both for the class of voting bodies formed by agents with complementary strategies and a complementary social decision rule, and for the even more general class of antisymmetric voting bodies. For a bird’s-eye view, the logical structure of the main symmetry results of this study is summarized in Fig. 4. Returning to two-component societies, we observe that in some cases, cooperating group members improve their position at the expense of individualists, which is disadvantageous for the whole society.
@@
Acknowledgments We are very grateful to two anonymous referees for their careful reading and especially for suggesting a generalization and the idea of reverse time to obtain a simpler proof of the key lemma. In the final version, we use a similar approach in terms of the antisymmetric voting body, which is necessary for further generalization of the result. We also thank the participants of the seminar on stochastic voting models at the Institute of Control Sciences of the Russian Academy of Sciences: Yana Tsodikova, Anton Loginov, Vitaly Malyshev, Vladislav Maksimov, Anna Khmelnitskaya, Irina Kalushina, Zoya Lezina, Vladimir Borzenko.